OpenStudy (anonymous):

Suppose that we know.. ∫1,10(3k(x)+5)dx=66 What is the value of.. ∫1,10k(x)dx

OpenStudy (anonymous):

What's your question?...

OpenStudy (anonymous):

Is 1,10 corresponding with a,b or b,a?

OpenStudy (anonymous):

a,b

OpenStudy (anonymous):

\[\int\limits_{1}^{10}(3kx+5)dx=\frac{ 3k }{ 2 }x^2+5x\] \[[\frac{ 3k }{ 2 }(10)^2+5(10)]-[\frac{ 3k }{ 2 }(1)^2+5(1)]=150k+50-\frac{ 3k }{ 2 }-5\] \[=\frac{297}{2}k+45\] \[\frac{297}{2}k+45=66\] \[k=\frac{14}{99}\] \[\int\limits_{1}^{10}(kx)dx=\frac{k}{2}x^2\] \[\frac{7}{99}(10)^2-\frac{7}{99}(1)^2=\frac{700}{99}-\frac{7}{99}\] \[=7\]

OpenStudy (anonymous):

I just started learning integrals. Can't believe I solved it for you...

OpenStudy (anonymous):

Thank you so much